# Scalable Formal Verification via Autoencoder Latent Space Abstraction

### v1

@@ 3.2. Notation
Given a high $n_{x}$-dimensional state $x$, an encoder $h_{enc}:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{p}}$ maps $x$ to an $n_{p}$-dimensional state $z=h_{enc}(x)$. We refer to $\mathbb{R}^{n_{p}}$ as the latent space. A decoder $h_{dec}:\mathbb{R}^{n_{p}}\rightarrow\mathbb{R}^{n_{x}}$ then maps a $z\in\mathbb{R}^{n_{p}}$ to a state $x$. Given a set of points $A$, $\text{Conv}(A)$ defines the convex hull of the set. In a metric space $\mathbb{R}^{n_{x}}$, $\mathcal{B}_{r}(x)=\{x^{\prime}\in\mathbb{R}^{n_{x}}\mid\|x-x^{\prime}\|\leq r\}$ denotes a ball of radius $r\geq 0$ centered at $x\in\mathbb{R}^{n_{x}}$.
@@ 4.3. Nondeterministic Transition System Abstraction
A Transition System (TS) is a tuple $\mathcal{N}=(Q,T,L,AP)$, where $Q$ is a finite set of states, $T:Q\times Q\rightarrow\{0,1\}$ is a transition function such that $T(q,q^{\prime})=1$ if $q^{\prime}$ is a successor state of state $q\in Q$, $AP$ is a set of atomic propositions, and $L:Q\rightarrow 2^{AP}$ is a labeling function that assigns to each $q\in Q$ a subset of $AP$. $\mathcal{N}$ is called Non-deterministic (NTS) if $\sum_{q^{\prime}\in Q}T(q,q^{\prime})>1$.
@@
To construct a finite NTS of System (1) in the latent space, we must first define the latent domain and its partitioning to obtain the set of discrete states $Q$. Next, we need a method to label regions of interest in the latent space that correspond to the regions of $R$ to define $L$ and $AP$, and a method to identify transitions between discrete states to construct $T$. We begin by discussing how regions in $\mathbb{R}^{n_{x}}$ are mapped into the latent space.
@@ 4.3.1. Mapping Regions of Interest
Given convex set $r_{i}\subset\mathbb{R}^{n_{x}}$, the encoding of this set $r_{Z,i}=\text{Post}_{h_{enc}}(r_{i})$ is also convex. Directly calculating $r_{Z,i}$ is generally infeasible; however, we can sample a set of $N$ points from $r_{i}$ and map each point through $h_{enc}$ to define a set in $\mathbb{R}^{n_{p}}$ where the convex hull is a subset of $r_{Z,i}$.
@@
We can formally bound $\epsilon(N,\delta)$ with $L_{h}\left(\frac{v_{n_{x}}}{Nv_{n_{p}}}\log{\frac{(4L_{h}\sqrt{n_{x}})^{n_{x}}}{\delta}}\right)^{1/{n_{p}}}$ where $v_{i}$ is the volume of an $i$-dimensional ball, as under the assumption that $r$ is defined as a subset of a unit ball and is uniformly sampled we can take $D(\partial r,\epsilon/2L_{h})\leq(4L_{h}\sqrt{n_{x}})^{n_{x}}$ and $\Lambda_{L_{h}}^{\epsilon}>\frac{v_{n_{p}}}{v_{n_{x}}L_{h}^{n_{p}}}\epsilon^{n_{p}}$ (lew2022simple). Note that $\epsilon$-RandUp does not require $h_{enc}$ to be a convex function or $r\in R$ to be a convex set, but the convexity of $h_{enc}$ enables us to identify both under- and over-approximations that are arbitrarily tight by increasing the number $N$ of samples used ((lew2022simple), Theorem 1). A visualization of the process is shown in Figure 2.
@@ 4.3.2. States and Labels
For a point $x\in X$, let $q_{x}\ni h_{enc}(x)$ denote the latent region in the partition $Q$ that contains $h_{enc}(x)$, and with an abuse of notation, let $L:X\to 2^{AP}$ or $L:Q\to 2^{AP}$, where $AP$ is defined in Eq. (8). Under the construction of the latent space described in Section 4.3.1 and labeling procedure described in Section 4.3.2, with high-confidence, it holds that $L(q_{x})\subseteq L(x)$.
Hence we capture (conservatively) all of the regions $R$ (and their negations) by the labels in $\mathcal{N}$, a detailed proof is available in Appendix A.3. We first translate $\varphi$ to NNF as $\mathit{NNF}(\varphi)$ (see Appendix A.1 for details), and then replace every $\neg p_{i}$ in $\mathit{NNF}(\varphi)$ with $p_{n,i}$. Then any trace that satisfies $\overline{\varphi}$ over $AP$ will also satisfy $\varphi$ over $\Pi$ (cauchi2019efficiency).
@@
4.4. NTS Transitions: Modeling the Latent Inclusion Dynamics
@@ A.2. Proof of Lemma 3 (connected inclusion dynamics)
where $L_{y}:=\{u\in\mathbb{R}^{t}:A_{2}(u)=y\}$ is an affine subspace of $\mathbb{R}^{t}$. Observe that $M$ is the continuous image of the affine space $TS$ under a coordinate-wise strictly monotone map, hence $M$ is a convex set defined with intervals over the first $n$ dimensions and constants for the next tnt-n dimensions. The intersection of convex sets is convex and hence connected. Therefore $f$ has connected pre-images.
@@ A.5. Proof of Theorem 11 (Sound Abstraction)
Assume $\omega_{x_{0}}\not\models\overline{\varphi}$, i.e. $\omega_{x_{0}}\models\neg\overline{\varphi}$. We begin with the base case $\overline{\varphi}:=p$. Then $\omega_{x_{0}}\not\models p$ means $p\notin L(x_{0})$ and since $L(q_{i})\subseteq L(x_{i})$ this implies $p\notin L(q_{0})$ and therefore $\omega_{q_{0}}\models\neg p$. Hence satisfaction of negative propositions in $AP$ is preserved from System (1) to $\mathcal{N}$. Since $\omega_{q_{0}}$ remains in relation with $\omega_{x_{0}}$, the satisfaction of negated propositions holds at subsequent states. By the semantics of LTL this satisfaction is applied inductively to Boolean and Temporal Operators. Then $\omega_{x_{0}}\not\models\overline{\varphi}$ implies $\exists\omega_{q_{0}}\not\models\overline{\varphi}$. This establishes the contrapositive and hence $\Omega_{q}\models\overline{\varphi}\implies\omega_{x}\models\overline{\varphi}$.

### v2

@@ 3.2. Notation
Given a high $n_{x}$-dimensional state $x$, an encoder $h_{enc}:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{p}}$ maps $x$ to an $n_{p}$-dimensional state $z=h_{enc}(x)$. We refer to $\mathbb{R}^{n_{p}}$ as the latent space. A decoder $h_{dec}:\mathbb{R}^{n_{p}}\rightarrow\mathbb{R}^{n_{x}}$ then maps a $z\in\mathbb{R}^{n_{p}}$ to a state $x$. Given a set of points AA, $\text{Conv}(A)$ defines the convex hull of the set. In a metric space $\mathbb{R}^{n_{x}}$, $\mathcal{B}_{r}(x)=\{x^{\prime}\in\mathbb{R}^{n_{x}}\mid\|x-x^{\prime}\|\leq r\}$ denotes a ball of radius $r\geq 0$ centered at $x\in\mathbb{R}^{n_{x}}$.
@@ 4.3. Nondeterministic Transition System Abstraction
A Transition System (TS) is a tuple $\mathcal{N}=(Q,T,L,AP)$, where $Q$ is a finite set of states, $T:Q\times Q\rightarrow\{0,1\}$ is a transition function such that $T(q,q^{\prime})=1$ if $q^{\prime}$ is a successor state of state $q\in Q$, APAP is a set of atomic propositions, and $L:Q\rightarrow 2^{AP}$ is a labeling function that assigns to each $q\in Q$ a subset of APAP. $\mathcal{N}$ is called Non-deterministic (NTS) if $\sum_{q^{\prime}\in Q}T(q,q^{\prime})>1$.
@@
To construct a finite NTS of System (1) in the latent space, we must first define the latent domain and its partitioning to obtain the set of discrete states $Q$. Next, we need a method to label regions of interest in the latent space that correspond to the regions of $R$ to define $L$ and APAP, and a method to identify transitions between discrete states to construct $T$. We begin by discussing how regions in $\mathbb{R}^{n_{x}}$ are mapped into the latent space.
@@ 4.3.1. Mapping Regions of Interest
Given convex set $r_{i}\subset\mathbb{R}^{n_{x}}$, the encoding of this set $r_{Z,i}=\text{Post}_{h_{enc}}(r_{i})$ is also convex. Directly calculating $r_{Z,i}$ is generally infeasible; however, we can sample a set of NN points from $r_{i}$ and map each point through $h_{enc}$ to define a set in $\mathbb{R}^{n_{p}}$ where the convex hull is a subset of $r_{Z,i}$.
@@
We can formally bound $\epsilon(N,\delta)$ with $L_{h}\left(\frac{v_{n_{x}}}{Nv_{n_{p}}}\log{\frac{(4L_{h}\sqrt{n_{x}})^{n_{x}}}{\delta}}\right)^{1/{n_{p}}}$ where $v_{i}$ is the volume of an $i$-dimensional ball, as under the assumption that $r$ is defined as a subset of a unit ball and is uniformly sampled we can take $D(\partial r,\epsilon/2L_{h})\leq(4L_{h}\sqrt{n_{x}})^{n_{x}}$ and $\Lambda_{L_{h}}^{\epsilon}>\frac{v_{n_{p}}}{v_{n_{x}}L_{h}^{n_{p}}}\epsilon^{n_{p}}$ (lew2022simple). Note that $\epsilon$-RandUp does not require $h_{enc}$ to be a convex function or $r\in R$ to be a convex set, but the convexity of $h_{enc}$ enables us to identify both under- and over-approximations that are arbitrarily tight by increasing the number NN of samples used ((lew2022simple), Theorem 1). A visualization of the process is shown in Figure 2.
@@ 4.3.2. States and Labels
For a point $x\in X$, let $q_{x}\ni h_{enc}(x)$ denote the latent region in the partition $Q$ that contains $h_{enc}(x)$, and with an abuse of notation, let $L:X\to 2^{AP}$ or $L:Q\to 2^{AP}$, where APAP is defined in Eq. (8). Under the construction of the latent space described in Section 4.3.1 and labeling procedure described in Section 4.3.2, with high-confidence, it holds that $L(q_{x})\subseteq L(x)$.
Hence we capture (conservatively) all of the regions $R$ (and their negations) by the labels in $\mathcal{N}$, a detailed proof is available in Appendix A.3. We first translate $\varphi$ to NNF as $\mathit{NNF}(\varphi)$ (see Appendix A.1 for details), and then replace every $\neg p_{i}$ in $\mathit{NNF}(\varphi)$ with $p_{n,i}$. Then any trace that satisfies $\overline{\varphi}$ over APAP will also satisfy $\varphi$ over $\Pi$ (cauchi2019efficiency).
@@
$4.$ NTS Transitions: Modeling the Latent Inclusion Dynamics
@@ A.2. Proof of Lemma 3 (connected inclusion dynamics)
where $L_{y}:=\{u\in\mathbb{R}^{t}:A_{2}(u)=y\}$ is an affine subspace of $\mathbb{R}^{t}$. Observe that $M$ is the continuous image of the affine space TSTS under a coordinate-wise strictly monotone map, hence $M$ is a convex set defined with intervals over the first $n$ dimensions and constants for the next tnt-n dimensions. The intersection of convex sets is convex and hence connected. Therefore $f$ has connected pre-images.
@@ A.5. Proof of Theorem 11 (Sound Abstraction)
Assume $\omega_{x_{0}}\not\models\overline{\varphi}$, i.e. $\omega_{x_{0}}\models\neg\overline{\varphi}$. We begin with the base case $\overline{\varphi}:=p$. Then $\omega_{x_{0}}\not\models p$ means $p\notin L(x_{0})$ and since $L(q_{i})\subseteq L(x_{i})$ this implies $p\notin L(q_{0})$ and therefore $\omega_{q_{0}}\models\neg p$. Hence satisfaction of negative propositions in APAP is preserved from System (1) to $\mathcal{N}$. Since $\omega_{q_{0}}$ remains in relation with $\omega_{x_{0}}$, the satisfaction of negated propositions holds at subsequent states. By the semantics of LTL this satisfaction is applied inductively to Boolean and Temporal Operators. Then $\omega_{x_{0}}\not\models\overline{\varphi}$ implies $\exists\omega_{q_{0}}\not\models\overline{\varphi}$. This establishes the contrapositive and hence $\Omega_{q}\models\overline{\varphi}\implies\omega_{x}\models\overline{\varphi}$.